OpenStudy (anonymous):

how do u find the slope intercept form of m= -3, (1, 3)?

OpenStudy (amistre64):

y = mx -mPx +Py

OpenStudy (amistre64):

where Px and Py come from P(x,y); P(1,3)

OpenStudy (anonymous):

the question on our packet says the formula is y=mx+b

OpenStudy (amistre64):

it is; and thats what it looks like when your done filling in the blanks

OpenStudy (amistre64):

y-Py = m(x-Px) y = mx -mPx + Py

OpenStudy (anonymous):

wheres p come ferom

OpenStudy (amistre64):

P is your point; P(x,y)

OpenStudy (anonymous):

whats the answer than

OpenStudy (amistre64):

fill it in and see what you get ... at least try

OpenStudy (amistre64):

m=-3; Px = 1; Py = 3 ... fill in the blanks